the number of number pairs (x,y) which will satisfy the equation x2-xy+y2=4(x+y-4)

Grade:Upto college level

3 Answers

nikhil tilak
14 years ago

by applying delta =0 condition we see that the above represents a pair of lines.so infinite points satisfy the equation.

Badiuddin askIITians.ismu Expert
14 years ago

Dear sudarshan

 x2-xy+y2=4(x+y-4)

  x2-x(y+4)+y2 -4y +16=0

if it has real solution then D≥0

so (y+4)2 -4( y2 -4y +16)≥0

 or (y-4)2  ≤0

 only possible when y=4

so for y=4 only one value of x=4 possible

so only one pair (4,4) is the solution

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
We are all IITians and here to help you in your IIT JEE & AIEEE preparation.

All the best.

Regards,
Askiitians Experts
Badiuddin

 

 

mycroft holmes
14 years ago

The equation is equivalent to (x-y)2+(x-4)2+(y-4)2 = 0 which means x=y=4 is the only solution.